# Towards Industrial-scale Product Configuration11footnotemark: 1

### v2

@@ 1 Introduction
The remainder of our paper is organized as follows: Section 2 introduces the CoomCore language and demonstrates its application through the KidsBike example. Section 3 presents our ASP-based approach to product configuration, detailing the refined fact format and ASP encoding. Section 4 outlines various extensions to the CoomCore language which contain commonly used features in product modeling and how they are handled within our ASP framework. These include concepts such as partonomy, cardinalities, and numerics, as well as user input and unbounded cardinalities. Further, an alternative encoding for solver flingo is presented. Finally, Section 5 describes the CoomSuite UI along with its ASP implementation through system clinguin and Section 6 demonstrates the practical value of the CoomSuite by evaluating alternative ASP solutions against benchmarks of varying complexity levels.
@@ 4.3 Solving Coom[x] with flingo
An alternative Coom[x] encoding is given for solver flingo (Cabalar et al. 2025)555fclingo in earlier versions. which is a prototype system for solving conditional linear constraints with integer variables in ASP. This system is a continuation of the lc2casp system (Cabalar et al. 2016) enhanced with conditional linear constraints given via the translation in (Cabalar et al. 2020). flingo uses CASP solver clingcon (Banbara et al. 2017) as a backend, therefore being able to deal with large integer ranges and numerical calculations. The main difference, however, is that variables in flingo can stay undefined while in clingcon all variables need to have a value assigned. As Coom allows for optional attributes, this is a crucial feature. The encoding for flingo is identical in many parts to the clingo encoding from the previous section, except for rules which treat integer attributes and numerical constraints.
We start by giving a brief overview of the flingo features we are utilizing for this adaptation and then proceed to highlight the differences between the two encodings. Our main objective with the flingo encoding is to improve performance for numerical instances by overcoming clingo’s grounding bottleneck. This is achieved in flingo by allowing for a special type of variables which are not subject to grounding. These are called integer variables and as their name suggests, can take values from the domain of the integers. Importantly, they are being treated differently than propositional variables. Further, to be able to properly replace the necessary rules from our earlier encoding, we require that these integer variables can be used to build linear constraints while comprising a notion of undefinedness and the possibility to define defaults. Lastly, we also need to be able to exclude variables from a calculation if they do not fulfill certain conditions, i.e., aggregate functions should allow for conditionality. flingo provides us with all these features as we see in the following. Rules in flingo are written in the same way as in clingo but additionally there are a few special theory atoms through which we can define integer variables and build up constraints over them. Here, we explain only those theory atoms that we are using in our encoding and refer the reader to A and Cabalar et al. (2025) for further documentation and examples. We proceed by showing snippets of the flingo encoding while highlighting the differences from the clingo encoding from Section 4.2.2.
@@ 7 Discussion
Appendix A flingo language
This section is taken and adapted from a forthcoming paper on the system flingo, and will be removed on publication.
We provide here a short introduction to the language of system flingo and illustrate its behavior with examples. The main feature of flingo is that the values assigned to integer variables need to be founded. Recall that in ASP, propositional atoms are false by default and any true propositional atom needs to be founded in the facts of the program by a sequence of rules. Analogously, in flingo, integer variables are undefined by default and any defined integer variable needs to be founded in the facts of the program by using the rules. This behavior is inspired by the Logic of Here-and-There with constraints ($\textrm{HT}_{\!c}$Cabalar et al. (2016)) and is significantly different from other existing CASP solvers such as clingcon. As an example, consider the program
{a}. &sum{x} = 1 :- a. (1)
which has exactly two flingo stable models: one where a is false and x is undefined, and another where a is true and x is assigned the value $1$. Both of these are also clingcon stable models, but clingcon has infinitely many stable models where a is false and x is assigned any integer. The second distinct feature of flingo is the support of aggregates over integer variables, something that is not supported by other existing CASP solvers.
We start by describing flingo’s conditional terms, which are inspired by aggregate elements in ASP and $\textrm{HT}_{\!c}$ conditional terms. A product term is either an integer, an integer variable, or an expression of the form nxn*x where $n$ is an integer and $x$ is an integer variable. A flingo conditional term is an expression of the form
$s:L$ (2)
where $s$ is a product term and $L$ is a comma separated list of propositional literals. The term $s$ is the one considered when the condition $L$ is satisfied, otherwise a default term is used, which is context-dependent and introduced below.
A flingo term is either a product term or a flingo conditional term. The following constraint atoms are supported by flingo:
${\texttt{\&sum}}\{t_{1};\ldots;t_{n}\}\prec s$ (3)
${\texttt{\&sus}}\{t_{1};\ldots;t_{n}\}\prec s$ (4)
${\texttt{\&max}}\{t_{1};\ldots;t_{n}\}\prec s$ (5)
${\texttt{\&min}}\{t_{1};\ldots;t_{n}\}\prec s$ (6)
${\texttt{\&df}}\{x\}$ (7)
where each $t_{i}$ is a flingo term, $s$ is an product term, $x$ is an integer variable, and $\prec$ is one of the following relations <=, =, !=, <, >, >=. The corresponding default terms are defined as 0 for $\mathtt{sum}$ and $\mathtt{sus}$, while $\mathtt{max}$ and $\mathtt{min}$ use the minimum and maximum elements of the integer domain, respectively. The operation $\mathtt{sus}$ stands for strict sum (more precisely $\mathtt{sus}$ stands for sum strict) while $\mathtt{sum}$ is its non-strict version. Intuitively, the non-strict sum discards undefined variables, while the strict does not. If any variable is undefined, the constraint atom with strict sum is false. As an example, consider the constraint atoms
${\texttt{\&sum}}\{x_{1};x_{2}\}\ \mbox{<=}\ 3$ (8)
${\texttt{\&sus}}\{x_{1};x_{2}\}\ \mbox{<=}\ 3$ (9)
and the valuation $v=\{x_{1}\mapsto 1\}$}. In this valuation, variable $x_{2}$ is undefined. Atom (8) is satisfied by valuation $v$ because $v(x_{1})=1\leq 3$ and $x_{2}$ is undefined, and thus is discarded from the sum. In contrast, atom (9) is not satisfied simply because $x_{2}$ is undefined. Operations $\mathtt{max}$ and $\mathtt{min}$ always work in a non-strict way: for instance, $\mathtt{max}$ always return the maximum value among the defined terms in the set.
A flingo atom is either a propositional atom or a constraint atom of the forms (3-7). A flingo literal is any flingo atom $b$, its default negation $\mathtt{not}\ b$ or a flingo atom preceded by two negations $\mathtt{not}\ \mathtt{not}\ b$. A flingo rule is an expression of the form
$a_{0}\ \texttt{:-}\ l_{1},\ldots,l_{n}$ (10)
where $a_{0}$ is a flingo atom or the symbol $\bot$ denoting falsity, and each $l_{i}$ is a flingo literal. Note that, syntactically, every clingcon rule is also a flingo rule.
One may raise the question of why we do not use &sum to denote the strict sum in flingo. The reason is that we aim to keep the syntax of flingo as close as possible to the usual ASP syntax for aggregates, and in ASP, aggregates behave in a non-strict way. For instance, the flingo program
a :- &sum{x : p} = 0. p.
has a single stable model where a and p are true and $x$ is undefined, while if we replace &sum with &sus, the resulting program leaves a undefined. This program can be rewritten in ASP using the aggregate #sum as follows:
a :- #sum{ X: val(X), p } = 0. p.
whose unique stable model is $\{\texttt{a}\}$.
The relation between flingo and clingcon is similar to the relation between the stable models semantics and classical propositional logic, except that here we need to restrict to stable models without undefined integer variables. Since flingo may leave integer variables undefined, a program in flingo may have stable models that the corresponding program in clingcon does not have. As an example, consider again the clingcon program consisting of the single rule $\texttt{a}\ \texttt{:-}\ {\texttt{\&sum}}\{x\}=x$} and its corresponding flingo program consisting of the single rule $\texttt{a}\ \texttt{:-}\ {\texttt{\&sus}}\{x\}=x$}. The first program has infinitely many clingcon answer sets where a is true and $x$ is assigned any integer, while the second program has a single flingo answer set where a is false and $x$ is undefined.
Continuing with the analogy between the stable models semantics and classical propositional logic, in the latter, we can capture all models of a propositional theory by adding choice rules for each propositional variable. Similarly, we can capture all stable models of a clingcon program by adding facts of the form ${\texttt{\&sum}}\{x\}=x$} for each integer variable $x$ in flingo. These facts act as a kind of choice rule for integer variables, allowing them to pick any value.
Assignments
A common pattern in flingo programs is to define a variable in terms of other variables. While constraint atoms (3-7) do not make any difference between the variables occurring on them, an assignment rule allow us to naturally express this pattern. We define an assignment rule as one of the two expressions
${\texttt{\&sus}}\{t_{1};\ldots;t_{n}\}\ \texttt{=:}\ s\ \ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m}$ (11)
${\texttt{\&in}}\{t_{l}..t_{u}\}\ \texttt{=:}\ s\ \ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m}$ (12)
where each $t_{i}$ and $t_{l}$, $t_{u}$ are flingo terms, $s$ is an product term, and each $l_{i}$ is a flingo literal. The operator $\mathtt{in}$ functions like a choice rule with a fixed range for integer variables, assigning a value in the range of $t_{l}$ and $t_{u}$ to $s$.
An assignment rule (11) or (12) only defines the variables in $s$ so that the constraint atom obtained by replacing the assignment operator =: with equality = is satisfied. However, it does not define any variables in $t_{1},\ldots,t_{n}$. For instance, the assignment rule &sus{x ; y} =: z ensures that z is defined as the sum of x and y, when both x and y are defined. However, if x or y are undefined, then the rule has no effect on the value assigned to z. As an example, if we add facts x = 1 and y = 2 to the above assignment, then the resulting program has a single stable model where x, y, and z are assigned $1$, 22, and $3$, respectively. If instead we add only the fact x = 1, then the resulting program has a single stable model where x is assigned $1$, while y and z are undefined. And if instead we add the facts x = 1 and z = 5, then the resulting program has a single stable model where x and z are assigned $1$ and $5$, respectively, while y is undefined.
An assignment rule of the form of (11) or (12) is an abbreviation, respectively, for the rules:
${\texttt{\&sus}}\{t_{1};\ldots;t_{n}\}\ \mbox{=}\ s\ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m},{\texttt{\&df}}(x_{1}),\dotsc,{\texttt{\&df}}(x_{k})$ (13)
${\texttt{\&in}}\{t_{l}..t_{u}\}\ \mbox{=}\ s\ \ \ \texttt{:-}\ \ \ l_{1},\ldots,l_{m},{\texttt{\&df}}(x_{1}),\dotsc,{\texttt{\&df}}(x_{k})$ (14)
where $x_{1},\ldots,x_{k}$ are all the the integer variables occurring in $t_{1},\dotsc,t_{n}$ or $t_{l}$ and $t_{u}$, respectively.

### v3

@@ 1 Introduction
The remainder of our paper is organized as follows: Section 2 introduces the CoomCore language and demonstrates its application through the KidsBike example. Section 3 presents our ASP-based approach to product configuration, detailing the refined fact format and ASP encoding. Section 4 outlines various extensions to the CoomCore language which contain commonly used features in product modeling and how they are handled within our ASP framework. These include concepts such as partonomy, cardinalities, and numerics, as well as user input and unbounded cardinalities. Further, an alternative encoding for solver flingo (Cabalar et al. 2026) is presented. Finally, Section 5 describes the CoomSuite UI along with its ASP implementation through system clinguin and Section 6 demonstrates the practical value of the CoomSuite by evaluating alternative ASP solutions against benchmarks of varying complexity levels.
@@ 4.3 Solving Coom[x] with flingo
An alternative Coom[x] encoding is given for solver flingo555fclingo in earlier versions. which is a prototype system for solving conditional linear constraints with integer variables in ASP. This system is a continuation of the lc2casp system (Cabalar et al. 2016) enhanced with conditional linear constraints given via the translation in (Cabalar et al. 2020). flingo uses CASP solver clingcon (Banbara et al. 2017) as a backend, therefore being able to deal with large integer ranges and numerical calculations. The main difference, however, is that variables in flingo can stay undefined while in clingcon all variables need to have a value assigned. As Coom allows for optional attributes, this is a crucial feature. The encoding for flingo is identical in many parts to the clingo encoding from the previous section, except for rules which treat integer attributes and numerical constraints.
We start by giving a brief overview of the flingo features we are utilizing for this adaptation and then proceed to highlight the differences between the two encodings. Our main objective with the flingo encoding is to improve performance for numerical instances by overcoming clingo’s grounding bottleneck. This is achieved in flingo by allowing for a special type of variables which are not subject to grounding. These are called integer variables and as their name suggests, can take values from the domain of the integers. Importantly, they are being treated differently than propositional variables. Further, to be able to properly replace the necessary rules from our earlier encoding, we require that these integer variables can be used to build linear constraints while comprising a notion of undefinedness and the possibility to define defaults. Lastly, we also need to be able to exclude variables from a calculation if they do not fulfill certain conditions, i.e., aggregate functions should allow for conditionality. flingo provides us with all these features as we see in the following. Rules in flingo are written in the same way as in clingo but additionally there are a few special theory atoms through which we can define integer variables and build up constraints over them. Here, we explain only those theory atoms that we are using in our encoding and refer the reader to Cabalar et al. (2026; 2025) for further documentation and examples. We proceed by showing snippets of the flingo encoding while highlighting the differences from the clingo encoding from Section 4.2.2.